Gucchi:

Object A has a mass of 12 g and a volume of 8 cm3. Object B has a mass of 20 g and a volume of 8 cm3. Which object has a greater density and by how much? Object A, by 1 gram per cubic centimeter Object A, by 8 grams per cubic centimeter Object B, by 1 gram per cubic centimeter Object B, by 8 grams per cubic centimeter

Gucchi:

am i right with C?

Gucchi:

@carmelle

carmelle:

remember, \[density =\frac{ mass }{ volume }\] so, how would you calculate the density of objects A and B? Would you like to do the math?

Gucchi:

12/8?

carmelle:

Object A: \[density =\frac{ 12 }{ 8 }=~?\] Object B: \[density =\frac{ 20 }{ 8 }=~?\] Once you figure that out, you subtract to find your final answer: Object `B` - Object `A` = ?

Gucchi:

1

Gucchi:

so C

carmelle:

Yup, thats correct, but I would recommend you work it out and also have this formula in ur notes because you might need it in the future :)

Gucchi:

ty